OpenStudy (vishweshshrimali5):

Find the last digit of 7 to the power 7 to the power 7 to the power 7 and so on upto 1 millionth term. Thanks in advance

OpenStudy (anonymous):

Why:)

OpenStudy (vishweshshrimali5):

Hey estudies please help me dude Its very urgent

OpenStudy (anonymous):

what ive noticed is that if you look at 7 and its powers, 7^1 ends in a 7, 7^2 ends in a 9, 7^3 ends in a 3, 7^4 ends in a 1, 7^5 ends in a 7...and so it repeats all the time...mayb that wud help

OpenStudy (anonymous):

there must be a patter yes? i \[7^7\] ends in a 3

OpenStudy (anonymous):

so i will be willing to bet they end up in a 3 from then on

OpenStudy (vishweshshrimali5):

Yes you are right but will you please explain step by step

OpenStudy (anonymous):

sorry i am thinking...

OpenStudy (anonymous):

ok lets try this explanation, although it may not be a very good one taking powers of 7, the remainders (when divided by ten) aka the ones digits, repeat in the pattern 1,9,3,1,9,3 ...

OpenStudy (anonymous):

nope i keep coming up with bad reasons still thinking

OpenStudy (vishweshshrimali5):

I can help you in this can i ?

OpenStudy (anonymous):

please. i mean it seems obvious by i cannot seem to get a grip

OpenStudy (vishweshshrimali5):

There is a phenomena called cyclicity

OpenStudy (vishweshshrimali5):

This can be applied here

OpenStudy (anonymous):

yes so it seems clear that if \[7^7\] ends in a 3 then so must \[7^{7^7}\] but i cannot come up with a good explanation.

OpenStudy (anonymous):

7^3k and 7^7k periodicity for 3 ending....

OpenStudy (anonymous):

oh yes now i can

OpenStudy (vishweshshrimali5):

This can be applied here

OpenStudy (vishweshshrimali5):

Thats good

OpenStudy (anonymous):

ok \[7^7\] ends in a 3 and therefore \[7^{7^7}\] is \[7^n\times 7^3\] where \[n\equiv 1(mod 3)\]

OpenStudy (anonymous):

no that is wrong

OpenStudy (anonymous):

this problem is not as trivial as i thought or else my mind is shot

OpenStudy (vishweshshrimali5):

Oh no that is wrong dude

OpenStudy (vishweshshrimali5):

Yaaa it might happen because this is an olympiad question

OpenStudy (anonymous):

I think it can't be 3^k, has to be 7^k by definition in the question.

OpenStudy (anonymous):

yeah i see that it is wrong

OpenStudy (anonymous):

7^1 = 7 7^2 = 49 7^3 = 343 7^4 = 2401 7^5 = 16807 7^6 = 117649 7^7 = 823543 last digit of 7^(3k and 7k) = 3 set k = 7^m

OpenStudy (vishweshshrimali5):

OKKKKKKKKKKKK Then ?

OpenStudy (anonymous):

It's just induction.